# TorchCP: A Python Library for Conformal Prediction

### v2

@@ A.3 Regression
ABS (Lei et al., 2018): The ABS score $V(\bm{x},y):=|f(\bm{x})-y|$ is designed to measure the absolute deviation of a prediction $f(\bm{x})$ from the observed target value $y$, where $f(\bm{x})$ is a point estimate of given $\{\bm{x}\}$. Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha},f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}].$
@@
A.4 Graph
Here, we consider node classification tasks. The Graph is represented as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}:=\{v_{i}\}_{i=1}^{N}$ denotes the node set and $\mathcal{E}$ denotes the edge set with $|\mathcal{E}|=E$. Let $\bm{X}:=[\bm{x}_{1},\cdot\cdot\cdot,\bm{x}_{N}]^{T}$ be the node feature matrix, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional feature vector for node $v_{i}$. The label of node $v_{i}$ is $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}:=\{1,2,...,K\}$ denotes the label space.
DAPS (Zargarbashi et al., 2023): DAPS proposes a diffusion-based method that incorporates neighborhood information by leveraging the network homophily. Specifically, the score function is defined as: $V(\bm{x}_{i},y,\lambda):=(1-\lambda)\hat{V}(\bm{x}_{i},y)+\frac{\lambda}{|\mathcal{N}_{i}|}\sum_{v_{j}\in\mathcal{N}_{i}}\hat{V}(\bm{x}_{j},y),$ where $\hat{V}(\cdot,\cdot)$ is the basic non-conformity score function, e.g., APS and RAPS, and $\lambda$ is a diffusion parameter.
@@ A.4 Graph
CF-GNN (Huang et al., 2023): CF-GNN introduces a topology-aware output correction model, akin to GCN, which employs a conformal-aware inefficiency loss to refine predictions and improve the efficiency of post-hoc CP. For the loss, CF-GNN adopts the optimizing efficiency version of ConfTr.
A.5 LLMs
@@
Conformal LLM (Quach et al., 2024): This paper proposes a novel conformal procedure based on risk control framework (Bates et al., 2021) for language models by reformulating the generation task as a sampling problem rather than enumerating all possible outputs. The core algorithm samples candidate outputs until a calibrated stopping rule is met, while using a rejection rule to filter out low-quality samples. The method guarantees that the output set contains at least one acceptable answer with high probability while maintaining the statistical guarantees of traditional conformal prediction.
Appendix B Benchmark Results
B.1 Classification
@@
In Figure 3, we present the experimental results of split CP on the ImageNet dataset, using THR, APS, RAPS, and SAPS. In Figure 4, we present a comparative analysis of different conformal prediction methods on the ImageNet dataset, including THR, APS, RAPS, Split CP, Class-wise CP, and Clustered CP. Clustered CP achieves the smallest average size and best class-conditional coverage.
@@
B.2 Regression
In Figure 5, we present an example of experimental results on the Community dataset (Redmond, 2002), using ABS, CQR, CQRR, CQRM, CQRFM, R2CCP with Split CP.
@@ B.2 Regression
B.3 Graph Node Classification
In Figure 6, we present an example of experimental results on the CoraML dataset (McCallum et al., 2000), using APS, DAPS, CF-GNN, and SNAPS. The experiments are conducted in a transductive learning setting for node classification tasks.
@@ B.3 Graph Node Classification
B.4 LLM
In Table 2, we present an example prediction set, generated by conformal LLM, of the TriviaQA dataset (Joshi et al., 2017). The results show that the generation process of the prediction set stops after sample 5, because the sum of scores collectively exceeds the conformal threshold.
Table 2: Example prediction set of conformal LLM on the TriviaQA dataset.
@@ B.4 LLM
Sample 1 Mutiny On The Bounty 0.5347 True
Sample 2 Treasure Island 0.5087 False
Sample 3 Treasure Island 0.5087 False
Sample 4 Mutiny on the Bounty 0.1889 True
Sample 5 Robinson Crusoe 0.1384 False
Appendix C Comparison of Computational Efficiency
We evaluated the computational efficiency of TorchCP by comparing its inference time with other existing libraries of conformal prediction. In particular, we focus on the time of calibration and inference on MNIST (LeCun et al., 2010) and ImageNet (Russakovsky et al., 2015) throughout the conformal prediction process. We leave out some libraries that cannot be implemented on the dataset without making significant changes. The results in Figure 7 show that TorchCP is the most efficient implementation among these libraries. TorchCP exhibits superior computational efficiency, attributed to its seamless integration with PyTorch’s architecture and mini-batch parallel processing capabilities.
@@ Appendix C Comparison of Computational Efficiency
Figure 7: Time comparison of conformal prediction using various open-sourced libraries on MNIST and ImageNet datasets.
Appendix D Extensibility
D.1 Add new scores
Adding new scores involves two steps:
  1. 1.
  2. 2.
D.2 Add new predictors
  1. 1.
    Implementing a New Predictor Class: new Predictors should be implemented in the predictor module, inheriting from the BasePredictor class. Using classification as an examples, one could alternatively choose to override the functions: __init__(), calibrate() and predict().
  2. 2.
D.3 Add new trainers
Adding new trainers is similar to steps of D.1 and D.2:
  1. 1.
  2. 2.
D.4 Add new metrics
Adding new metrics involves two steps:
  1. 1.
  2. 2.
    Adding an Interface: After customizing the metric function, register it by Registry("METRICS") in the utils/metrics.py.

### v3

@@ A.3 Regression
ABS (Papadopoulos et al., 2002): The ABS score $V(\bm{x},y):=|f(\bm{x})-y|$ is designed to measure the absolute deviation of a prediction $f(\bm{x})$ from the observed target value $y$, where $f(\bm{x})$ is a point estimate of given $\{\bm{x}\}$. Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha},f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}].$
NORABS (Vovk et al., 2005): The NORABS score $V(\bm{x},y):=\frac{|f(\bm{x})-y|}{\hat{\sigma}(\bm{x})}$ is a normalized variant of ABS. It uses models $f(\bm{x})$ and $\hat{\sigma}(\bm{x})$ trained to maximize the likelihood of the data with respect to $\mathbb{E}[Y_{\text{test}}|X_{\text{test}}=\bm{x}]$ and $\sqrt{\text{Var}[Y_{\text{test}}|X_{\text{test}}=\bm{x}]}$, respectively. Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha}\cdot\hat{\sigma}(\bm{x}_{n+1}),f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}\cdot\hat{\sigma}(\bm{x}_{n+1})].$
@@
AgACI (Zaffran et al., 2022): AgACI is a parameter-free conformal prediction method for time series that builds on ACI by using online expert aggregation. It runs multiple models in parallel with different learning rates $\gamma_{k}$ and adaptively aggregates their prediction intervals using weighted averages, where weights are based on the historical pinball loss. This approach eliminates the need to tune $\gamma$ manually and ensures valid coverage while improving interval efficiency, especially under temporal dependencies.
A.4 GNN
We consider the prediction tasks on graph data. The Graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}:=\{v_{i}\}_{i=1}^{N}$ denotes the node set and $\mathcal{E}$ denotes the edge set with $|\mathcal{E}|=E$. Let $\bm{X}:=[\bm{x}_{1},\cdot\cdot\cdot,\bm{x}_{N}]^{T}$ be the node feature matrix, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional feature vector for node $v_{i}$. The label of $v_{i}$ is $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}:=\{1,2,...,K\}$ denotes the label space.
DAPS (Zargarbashi et al., 2023): DAPS proposes a diffusion-based method that incorporates neighborhood information by leveraging the network homophily. Specifically, the score function is defined as: $V(\bm{x}_{i},y,\lambda):=(1-\lambda)\hat{V}(\bm{x}_{i},y)+\frac{\lambda}{|\mathcal{N}_{i}|}\sum_{v_{j}\in\mathcal{N}_{i}}\hat{V}(\bm{x}_{j},y),$ where $\hat{V}(\cdot,\cdot)$ is the basic non-conformity score function, e.g., APS, and $\lambda$ is a diffusion parameter.
@@ A.4 GNN
CF-GNN (Huang et al., 2023): CF-GNN introduces a topology-aware output correction model, akin to Graph Convolutional Network, which employs a conformal-aware inefficiency loss to refine predictions and improve the efficiency of post-hoc CP. For the loss, CF-GNN adopts the optimizing efficiency version of ConfTr.
A.5 LLM
@@
Conformal LLM (Quach et al., 2024): This paper proposes a novel conformal procedure based on a risk control framework (Bates et al., 2021) for language models by reformulating the generation task as a sampling problem rather than enumerating all possible outputs. The core algorithm samples candidate outputs until a calibrated stopping rule is met, while using a rejection rule to filter out low-quality samples. The method guarantees that the output set contains at least one acceptable answer with high probability while maintaining the statistical guarantees of traditional conformal prediction.
A.6 Metrics
In this section, we provide a brief overview of the evaluation metrics supported by TorchCP for various predictive tasks, as summarized in Table 4. These metrics facilitate rigorous and fine-grained assessment of prediction sets, offering deep insights into model uncertainty.
Table 4: Implemented evaluation metrics in TorchCP.
Task Metrics
Classification Coverage, Size, CovGap (Ding et al., 2024), WSC (Romano et al., 2020), SSCV (Angelopoulos et al., 2020), Violated classes (Kasa and Taylor, 2023), DiffViolation (Angelopoulos et al., 2020), , etc.
Regression Coverage, Size
GNN Coverage, Size, Singleton ratio (Zargarbashi et al., 2023)
LLM SetLoss, SSCL (Quach et al., 2024)
Appendix B Extensibility
B.1 Add new scores
Adding new scores involves two steps:
  1. 1.
  2. 2.
B.2 Add new predictors
  1. 1.
    Implementing a New Predictor Class: new Predictors should be implemented in the predictor module, inheriting from the BasePredictor class. Using classification as an example, one could alternatively choose to override the functions: __init__(), calibrate(), and predict().
  2. 2.
B.3 Add new trainers
Adding new trainers is similar to steps of B.1 and B.2:
  1. 1.
  2. 2.
B.4 Add new metrics
Adding new metrics involves two steps:
  1. 1.
  2. 2.
    Adding an Interface: After customizing the metric function, register it by Registry(‘‘METRICS’’) in the utils/metrics.py.
Appendix C Example Codes for TorchCP
TorchCP provides a collection of example codes in the “examples" folder  [^6] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples of the official codebase, designed to help users quickly get started with various conformal prediction tasks. In this section, we briefly introduce representative examples for the four supported tasks.
C.1 Example codes for offline conformal prediction
This part provides brief descriptions of example codes under the offline setting:
  1. 1.
    Classification. This example code [^7] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100.py applies Split CP with the LAC score on CIFAR-100 using a pretrained ResNet-20. It evaluates the coverage and average set size of the resulting prediction sets.
  2. 2.
    Regression. The example code of regression [^8] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/regression_cqr_synthetic.py shows the complete workflow of applying Split CP with the Conformalized Quantile Regression score (Romano et al., 2019) on a synthetic dataset in regression task, including model training, calibration, prediction interval generation, and evaluation.
  3. 3.
    GNN. The node classification example [^9] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/gnn_transductive_coraml.py implements a Graph Convolutional Network (GCN) (Kipf and Welling, 2017) for node classification on the CoraML dataset (McCallum et al., 2000) under a transductive setting. Conformal prediction is then applied to generate prediction sets with guaranteed coverage.
  4. 4.
    LLM. The LLM example [^10] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/llm_ConformalLM_TriviaQA.py demonstrates how to deploy conformal prediction in large language model generation tasks, i.e., TriviaQA (Joshi et al., 2017).
C.2 Example codes for online conformal prediction
Online conformal prediction means that after predicting a test point, its true label is immediately revealed, and the test sample is incorporated into the calibration set. TorchCP supports this setting as well, with the following examples:
  1. 1.
    Regression. In TorchCP, the online setup of the regression task is integrated into ACIPredictor (Gibbs and Candes, 2021). Therefore, we added an example code  [^11] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/timeseries_aci_synthetic.py for the regression task in TorchCP.
  2. 2.
    Classification. For classification tasks, this online calibration approach can be naturally implemented in TorchCP by iteratively invoking the “calibrate()” function with new incoming data. To clarify this, we have added an online conformal prediction example 121212 https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100_online.py for a classification task in TorchCP.
Appendix D Empirical Results
This section presents empirical results of TorchCP. We first reproduce several classical conformal prediction methods across various tasks. Then, we benchmark the computational efficiency of TorchCP against several representative conformal prediction libraries.
D.1 Evaluation results
Classification.
@@ D.1 Evaluation results
In Figure 2, we present the experimental results of Split CP on ImageNet, using LAC, APS, RAPS (penalty=0.1, kreg =0), and SAPS (weight=0.2). In Figure 3, we present a comparative analysis of different conformal prediction methods on the ImageNet dataset, including Split CP, Class-conditional CP, and Clustered CP with APS. The number of trials is 1000.
@@
Figure 3: Conditional coverage results on ImageNet image classification with $\alpha=0.1$.
Regression.
In Figure 4, we present an example of experimental results on the Community dataset (Redmond, 2002), using ABS, CQR, CQRR, CQRM, CQRFM, R2CCP with Split CP. The number of trials is 1000.
@@
GNN.
In Figure 5, we present experimental results on the CoraML dataset (McCallum et al., 2000) for node classification under a transductive learning setting. We evaluate four methods: APS, DAPS, CF-GNN, and SNAPS, each conducted over 1000 independent trials. For DAPS, the diffusion parameter is set to 0.5. In CF-GNN, the topology-aware correction model is implemented using a two-layer GCN. For SNAPS, the hyperparameters are set as follows: $\lambda=\frac{1}{3}$, $\mu=\frac{1}{3}$, and $k=20$. These hyperparameter choices follow the default or recommended configurations reported in the respective original papers.
@@
LLM.
In Table 5, we present an example prediction set, generated by conformal LLM, of the TriviaQA dataset (Joshi et al., 2017). The generation process terminates after the fifth sample, as the cumulative score surpasses the conformal threshold.
Table 5: An Example of prediction sets from Conformal LLM on TriviaQA.
@@
Answer 1 Mutiny On The Bounty 0.5347 True
Answer 2 Treasure Island 0.5087 False
Answer 3 Treasure Island 0.5087 False
Answer 4 Mutiny on the Bounty 0.1889 True
Answer 5 Robinson Crusoe 0.1384 False
D.2 Comparison of computational efficiency
D.2.1 Classification Task
We evaluated the computational efficiency of TorchCP in image classification tasks by comparing its calibration and inference time with several existing CP libraries. Specifically, we use two standard benchmarks: MNIST (LeCun et al., 2010) and ImageNet (Russakovsky et al., 2015). Throughout all experiments, we used Split CP with APS as the conformal prediction method. Moreover, TorchCP supports two data processing modes: full-batch (processing the entire dataset at once), and batch (processing data in mini-batches). In our experiments, we set the batch size to 128 for batch processing. We use NVIDIA H100 for GPU computation and AMD EPYC 9654 for CPU computation.
MNIST. We compared the computational efficiency of five libraries: PUNCC, MAPIE, Fortuna, Crepes, and TorchCP (GPU-based full-batch processing). In the dataset, we use 48,000 samples for training, 6,000 for calibration, and 6,000 for testing. For TorchCP, we use a LeNet-5 model for training and inference using PyTorch on a GPU. For PUNCC, MAPIE, and Crepes, we obtain model predictions using PyTorch on a GPU, and then convert output probabilities to NumPy arrays for conformal calibration and inference using CPU. For Fortuna, we train the LeNet-5 model using the built-in training algorithm and perform model predictions with JAX on a GPU. We only report the total runtime for Fortuna as it ties the conformal calibration and inference processes together. Figure 6(a) presents the running times of the five libraries on the MINST dataset. The results show that TorchCP achieves the lowest calibration and inference time among the five libraries. Specifically, TorchCP completes calibration and inference for the entire dataset in just 0.07 seconds, using only 3.6% of the runtime needed by PUNCC (1.93 seconds). Moreover, the Crepes library exhibits the longest inference time on the MNIST dataset, likely due to its computation of p-values for each test sample’s non-conformity scores by comparing them against the entire calibration set. This substantial improvement highlights the efficiency advantage of TorchCP on small-scale datasets.
ImageNet. We compared the computational efficiency of four libraries: PUNCC, MAPIE, Crepes, and TorchCP (GPU-based batch processing). In the dataset, we use a pre-trained ResNet-50 model from torchvision and select 25,000 images from the validation set for calibration and another 25,000 for testing. For TorchCP, we leverage GPU-based batch processing to efficiently run the conformal pipeline. For PUNCC, MAPIE, and Crepes, we conducted experiments on the CPU, as they lack batch processing support, and their GPU-based full-batch processing causes out-of-memory errors. Figure 6(b) shows the results of running times of three libraries on the ImageNet dataset. The results demonstrate that TorchCP achieves the shortest calibration and inference times, cutting approximately 90% of the time required by PUNCC (104s vs. 995s). This significant enhancement underscores the scalability of TorchCP on large-scale datasets and models.
@@ D.2.1 Classification Task
Figure 6: Runtime comparison of different conformal prediction libraries.
Choices of Batch processing. To assess the efficiency of two processing modes (Full-batch and batch), we conduct experiments with TorchCP on both MNIST and ImageNet, using the same data splits and models as in prior experiments. We utilize GPU-accelerated computing for all experiments except the full-batch mode on ImageNet, where the dataset’s size exceeds GPU memory capacity. Figures 7(a) and  7(b) present the running times of different data processing modes on MNIST and ImageNet, respectively. On the small-scale dataset - MNIST, using full-batch processing on GPU achieves the fastest implementation among the three choices, requiring only 26% computational time of that with CPU. When it comes to the ImageNet dataset, GPU-based batch processing completes calibration and inference in approximately 104.29 seconds, compared to 1022.55 seconds for CPU-based processing, achieving a 9.8× speedup. These findings suggest that GPU-based full-batch processing is optimal for small-scale datasets where GPU memory is sufficient to load the whole dataset. Conversely, GPU-based batch processing is preferable for large-scale datasets like ImageNet, offering significant runtime advantages and enabling scalable, efficient deployment.
(a) MNIST
(b) ImageNet
Figure 7: Runtime comparison of different data processing modes.
D.2.2 Regression Task
We evaluate the computational efficiency and predictive performance of TorchCP and an R-based package, AdaptiveConformal (Susmann et al., 2023) on a time series regression task, using the France electricity forecasting dataset [^13] https://cran.r-project.org/web/packages/opera/index.html. Specifically, we use the first 350 time steps to train a SARIMA model with order (1, 0, 0) and seasonal order (1, 1, 0), and construct prediction intervals for the remaining time steps. To assess computational efficiency, we apply the ACI algorithm(Gibbs and Candes, 2021) using both TorchCP and AdaptiveConformal. Both AdaptiveConformal and TorchCP are executed on CPU in this experiment to ensure a fair comparison. Moreover, since ACI performs recalibration for each test instance, the reported time includes both calibration and inference.
Figure 8(a) reports the runtime results of these two packages. The results show that TorchCP achieves significantly higher computational efficiency compared to the R-based package, reducing running time from 6.1 seconds to around 0.06 seconds. Figures 8(b) and 8(c) illustrate the prediction intervals produced by AdaptiveConformal and TorchCP, respectively. The two sets of intervals are nearly identical, demonstrating the correctness and consistency of the TorchCP implementation. This substantial improvement highlights the efficiency advantage of TorchCP for time series regression.
(a) Time
(b) AdaptiveConformal
(c) TorchCP
Figure 8: Comparison of time consumption and prediction intervals between AdaptiveConformal and TorchCP on the France electricity dataset.
Appendix E The Advantage of our modular design
The modular design of TorchCP prioritizes low coupling to ensure flexibility in integrating diverse training algorithms, non-conformity score functions, and prediction workflows, accommodating the wide variety of CP methods and tasks. In contrast, previous packages (e.g., PUNCC and Fortuna) focus on a monolithic design where predictors and scoring functions are often predefined within a specific predictor (e.g., Split CP). This limits extensibility when new CP predictors or custom score functions are needed. TorchCP ’s decoupled design allows users to independently select or define new trainers, scores, or predictors, enabling flexible integration of different components. In what follows, we elaborate on the concrete benefits of our modular design:
  • Flexible combination: users can freely combine Trainer, Score, and Predictor modules to customize conformal prediction workflows. For instance, a practitioner can use different non-conformity scores (e.g., LAC or APS) within any predictor or trainer, while other packages only implement scores for predictors.
  • Modification isolation: our module design ensures that changes to one module (e.g., refining the APS score) do not affect other modules. This isolation minimizes the risk of unintended side effects, reducing maintenance complexity and preventing errors from propagating across the system. For instance, tightly coupled systems like PUNCC require redefining the predictor to modify a scoring function, while TorchCP confines changes to the target module, streamlining updates and preserving stability.
  • Research extensibility: by adhering to standardized interfaces, new functionality (e.g., a novel non-conformity score) can be integrated into TorchCP without altering existing modules. This contrasts with PUNCC, where introducing new scoring methods demands extensive predictor modifications. Extensibility supports rapid iteration and long-term evolvability, enabling efficient adaptation to new research or applications.
  • Ease of testing and debugging: decoupled modules can be tested and debugged independently, as their interactions are governed by well-defined interfaces. This modularity simplifies identifying and resolving issues within a single module (e.g., debugging a Trainer algorithm) without navigating complex dependencies. In TorchCP, this reduces testing overhead and accelerates development cycles compared to tightly coupled systems, where debugging often requires analyzing the entire pipeline.